# Two-particle calculations with quantics tensor trains - solving the parquet equations

### v1

@@ A.1 Parquet picture
$F(1,2,3,4)$ $=\Lambda(1,2,3,4)+\Phi^{ph}(1,2,3,4)$
$+\Phi^{\overline{ph}}(1,2,3,4)+\Phi^{pp}(1,2,3,4).$ (29)
@@
$F(1,2,3,4)$ $=\Gamma^{ph}(1,2,3,4)+\Phi^{ph}(1,2,3,4),$
$\Phi^{ph}(1,2,3,4)$ $=\Gamma^{ph}(1,2,5,6)G(6,7)G(8,5)F(7,8,3,4),$ (32a)
$F(1,2,3,4)$ $=\Gamma^{pp}(1,2,3,4)+\Phi^{pp}(1,2,3,4),$
$\Phi^{pp}(1,2,3,4)$ $=\frac{1}{2}\Gamma^{pp}(1,5,3,6)G(6,7)G(5,8)F(7,2,4,8),$ (32b)
where the summation over repeated arguments ($5,6,7$, and $8$) is implied and we also used a two-frequency two-spin notation for the one-particle Green’s function $G(1,2)=G_{\sigma_{1}\sigma_{2}}(\nu_{1},\nu_{2})$. From this representation it is visible what the true difference between the ph and pp BSEs is – the Green’s functions connect the vertices differently, i.e., different frequency arguments are summed over. Practical evaluation of these equations requires however the introduction of two different three-frequency (two fermionic, one bosonic) parametrizations. These parametrizations are sometimes called channel native.
A.2 Bethe-Salpeter picture / Channel native description
Following Ref. 35, in this work we use the following frequency convention for the ph and pp channels
@@ A.2 Bethe-Salpeter picture / Channel native description
Applying the above parametrizations to Eqs. (32) and additionally introducing the $d/m/s/t$ spin combinations leads to Eqs. (12). Now we however need channel transformations (frequency shifts) to evaluate the parquet equation. We need $F$ both in the ph and pp notations for the $d/m$ and $s/t$ channels, respectively. The channel transformations can be derived by going back and forth from three frequency to four frequency representations, e.g.
$F_{pp}(\nu,\nu^{\prime},\omega)=F_{pp}(-\nu_{1},\nu_{4},\nu_{1}+\nu_{2})$
$=F_{ph}(-\nu_{1},\nu_{4},-\nu_{2}-\nu_{4})=F_{ph}(\nu,\nu^{\prime},-\omega-\nu-\nu^{\prime}),$ (34)
@@
$p$ $h$$\longrightarrow$ $\quad pp$ (35)
$(\nu,\nu^{\prime}$ $,\omega)$$\longrightarrow$ $\quad(\nu,\nu^{\prime},-\omega-\nu-\nu^{\prime}).$
To use the crossing symmetry relation (31), we also need the ph to $\overline{ph}$ channel transformation. The crossing transformation means exchanging either first with third or second with fourth frequency, so we can write
$F_{\overline{ph}}(\nu,\nu^{\prime},\omega)=F_{\overline{ph}}(-\nu_{1},\nu_{4},\nu_{1}+\nu_{2})=$
$=F_{ph}(-\nu_{1},\nu_{2},\nu_{1}+\nu_{4})=F_{ph}(\nu,\nu+\omega,\nu^{\prime}-\nu)$ (36)
@@
$p$ $h$$\longrightarrow$ $\quad\overline{ph}$ (37)
$(\nu,\nu^{\prime}$ $,\omega)$$\longrightarrow$ $\quad(\nu,\nu+\omega,\nu^{\prime}-\nu).$
@@ Appendix B Affine transformations
Figure 13: MPO for affine transformations applied to a QTT. The leftmost virtual bonds of the MPO are contracted with a vector $\mathbf{1}=(1,\ldots,1)^{T}$ and the rightmost virtual bonds are projected to the carry of 0.
Let $\bm{y}$ and $\bm{x}$ be vectors with $M$ and $N$ integer components, respectively. An affine transformation is a map $\bm{x}\mapsto\bm{y}$ satisfying:
$\bm{y}=\tfrac{1}{s}(A\bm{x}+\bm{b}),$ (38)
where $s$ is integer, $A$ is a $M\times N$ matrix of integers, and $\bm{b}$ is an integer vector with $M$ components. Equivalently, we can set $s=1$ and allow $A$ and $\bm{b}$ to have entries in $\mathbb{Q}$.
Given a function $g(\bm{y})$, we construct a new function $f(\bm{x})$ by a coordinate transformation:
$f(\bm{x}):=g(\bm{y}(\bm{x}))$ (39)
whenever there exists $\bm{y}$ that satisfies Eq. (38), and $f(\bm{x})=0$ otherwise. We call this type of transformation a passive affine transformation, where for a given $\bm{x}$, we define the value of the new function $f(\bm{x})$ by picking the value of the old function $g(\bm{x})$ at the transformed point $\bm{y}$.
By introducing the residuum
$\bm{z}(\bm{y},\bm{x}):=A\bm{x}+\bm{b}-s\bm{y}$ (40)
and the corresponding transformation function:
$T(\bm{y},\bm{x}):=\begin{cases}1&\bm{z}(\bm{y},\bm{x})=0,\\ 0&\mbox{else},\end{cases}$ (41)
we can rewrite Eq. (39) as a tensor product:
$f(\bm{x})=\sum_{\bm{y}}T(\bm{x},\bm{y})g(\bm{y}),$ (42)
where the “inverse” transformation appears. This inverse transformation can again be written in the form (38), since the inverse of a nonsingular rational matrix is again rational.
In order to construct an MPO for Eq. (41), it is useful to start with “fused” rather than interleaved indices, i.e., we only separate out the length scales of the vector, but not its components:
@@
where $\bm{x}_{r}\in\{0,1\}^{N}$. (Thus, the legs $\bm{x}_{r}$ of the corresponding MPS are of dimension $2^{N}$ rather than $2$.) We perform similar decompositions for $\bm{y}$, $\bm{z}$, and $\bm{b}$. Consequentially, $T$ is decomposed as the following MPO:
$T(\bm{x},\bm{y})=[T_{1}]^{\bm{x}_{1}\bm{y}_{1}}_{\sigma_{1}}[T_{2}]^{\bm{x}_{2}\bm{y}_{2}}_{\sigma_{1}\sigma_{2}}\cdots[T_{R}]^{\bm{x}_{R}\bm{y}_{R}}_{\sigma_{R-1}},$ (44)
where again summation over duplicate indices is implied (cf. Sec. IV).
Let us first assume $s$ to be an odd number. Then Eq. (40) implies for the finest scale $r=R$:
$2\bm{c}_{R}+\bm{z}_{R}=A\bm{x}_{R}+\bm{b}_{R}-s\bm{y}_{R},$ (45)
where $\bm{c}_{R}$ is a vector of integers not necessarily confined to 0 and 1. $\bm{z}_{R}$ must vanish, which translates to the following constraint:
$\bm{y}_{R}\equiv A\bm{x}_{R}+\bm{b}_{R}\mod 2.$ (46)
@@
$\bm{c}_{R}$ is the outgoing carry of the calculation, which we must transmit to the previous core tensor $T_{R-1}$. To do so, we collect all distinct values of the carry for all possible inputs $\bm{x}_{R}$ into a tuple $(\bm{c}_{R1},\ldots,\bm{c}_{RD_{R-1}})$. The core tensor is then given by:
$[T_{R}]^{\bm{x}_{R}\bm{y}_{R}}_{\sigma}=\begin{cases}1&2\bm{c}_{R\sigma}=A\bm{x}_{R}+\bm{b}_{R}-s\bm{y}_{R},\\ 0&\mbox{else}.\end{cases}$ (47)
For all the other scales $r$, we must take into account the incoming carry $\bm{c}_{r+1}$ and must thus amend Eq. (45) to:
$2\bm{c}_{r}+\bm{z}_{r}=A\bm{x}_{r}+\bm{b}_{r}-s\bm{y}_{r}+\bm{c}_{r+1},$ (48)
@@
which implies, assuming again that we have collected all distinct outgoing carry values for all possible $\bm{x}_{r}$ and $\bm{c}_{r+1}$ into $(\bm{c}_{r\sigma})_{\sigma=1,\ldots,D_{r-1}}$, the following expression for the core tensor:
@@
$[T_{r}]^{\bm{x}_{r}\bm{y}_{r}}_{\sigma\sigma^{\prime}}=\begin{cases}1&2\bm{c}_{r\sigma}=A\bm{x}_{r}+\bm{b}_{r}-s\bm{y}_{r}+\bm{c}_{r+1,\sigma^{\prime}},\\ 0&\mbox{else}.\end{cases}$ (49)
The bond dimension of the tensors constructed thusly is likely optimal.
Different boundary conditions translate to different strategies of handling the leftmost carry $\bm{c}_{1}$: hard boundary conditions can be enforced by discarding all elements from $T_{1}$ except $c_{1}=0$. Periodic boundary conditions are implemented by summing $T_{1}$ over all possible values for $c_{1}$, or, equivalently, contracting $\sigma$ in Eq. (49) with a vector of ones, $\bm{1}=(1,1,\ldots,1)^{T}$. The resulting MPO is schematically depicted in Fig. 13.
If $2^{-k}s$ is odd, the construction is almost identical, but instead of $x_{r}$ and $y_{r}$ appearing in $T_{r}$, we have to shift the $y$ indices by $k$ places, i.e., we have that:
$[T_{r}]^{\bm{x}_{r}\bm{y}_{r+k}}_{\sigma\sigma^{\prime}}=\begin{cases}1&2\bm{c}_{r\sigma}=A\bm{x}_{r}+\bm{b}_{r}-\frac{s}{2^{k}}\bm{y}_{r+k}+\bm{c}_{r+1,\sigma^{\prime}},\\ 0&\mbox{else}.\end{cases}$ (50)
Once the MPO is constructed in this way, we fuse the $M$ virtual bonds into a single virtual bond, and then transform the MPO to the interleaved representation by splitting the MPO tensors using a QR decomposition.

### v2

@@ A.1 Parquet picture
$F(1,2,3,4)$ $=\Lambda(1,2,3,4)+\Phi^{\textit{ph}}(1,2,3,4)$
$\quad+\Phi^{\overline{\textit{ph}}}(1,2,3,4)+\Phi^{\textit{pp}}(1,2,3,4).$ (27)
@@
$F(1,2,3,4)$ $=\Gamma^{\textit{ph}}(1,2,3,4)+\Phi^{\textit{ph}}(1,2,3,4),$ (30a)
$\Phi^{\textit{ph}}(1,2,3,4)$ $=\Gamma^{\textit{ph}}(1,2,5,6)G(6,7)G(8,5)F(7,8,3,4),$
$F(1,2,3,4)$ $=\Gamma^{\textit{pp}}(1,2,3,4)+\Phi^{\textit{pp}}(1,2,3,4),$ (30b)
$\Phi^{\textit{pp}}(1,2,3,4)$ $=\tfrac{1}{2}\Gamma^{\textit{pp}}(1,5,3,6)G(6,7)G(5,8)F(7,2,4,8),$
where the summation over repeated arguments ($5,6,7$, and $8$) is implied and we also used a two-frequency two-spin notation for the one-particle Green’s function $G(1,2)=G_{\sigma_{1}\sigma_{2}}(\nu_{1},\nu_{2})$. This representation reveals the true difference between the ph and pp BSEs—the Green’s functions connect the vertices differently, i.e., different frequency arguments are summed over. Practical evaluation of these equations requires however the introduction of two different three-frequency (two fermionic, one bosonic) parametrizations. These parametrizations are sometimes called channel native.
A.2 Bethe–Salpeter picture / Channel native description
Following Ref. 37, in this work we use the following frequency convention for the ph and pp channels
@@ A.2 Bethe–Salpeter picture / Channel native description
Applying the above parametrizations to Eqs. (30) and additionally introducing the $d/m/s/t$ spin combinations leads to Eqs. (12). Now, however, we need channel transformations (frequency shifts) to evaluate the parquet equation. We need $F$ both in the ph and pp notations for the $d/m$ and $s/t$ channels, respectively. The channel transformations can be derived by going back and forth from three frequency to four frequency representations, e.g.
$F_{\textit{pp}}(\nu,\nu^{\prime},\omega)=F_{\textit{pp}}(-\nu_{1},\nu_{4},\nu_{1}+\nu_{2})$ (32)
$\quad=F_{\textit{ph}}(-\nu_{1},\nu_{4},-\nu_{2}-\nu_{4})=F_{\textit{ph}}(\nu,\nu^{\prime},-\omega-\nu-\nu^{\prime}),$
@@
ph $\longrightarrow\textit{pp}$ (33)
$(\nu,\nu^{\prime},\omega)$ $\longrightarrow(\nu,\nu^{\prime},-\omega-\nu-\nu^{\prime}).$
To use the crossing symmetry relation (29), we also need the ph to $\overline{\textit{ph}}$ channel transformation. The crossing transformation means exchanging either first and third or second and fourth frequency, so we can write
$F_{\overline{\textit{ph}}}(\nu,\nu^{\prime},\omega)=F_{\overline{\textit{ph}}}(-\nu_{1},\nu_{4},\nu_{1}+\nu_{2})=$ (34)
$\quad=F_{\textit{ph}}(-\nu_{1},\nu_{2},\nu_{1}+\nu_{4})=F_{\textit{ph}}(\nu,\nu+\omega,\nu^{\prime}-\nu)$
@@
ph $\longrightarrow\overline{\textit{ph}}$ (35)
$(\nu,\nu^{\prime},\omega)$ $\longrightarrow(\nu,\nu+\omega,\nu^{\prime}-\nu).$
@@ Appendix B Affine transformations
An important subset of transformations on a QTT are coordinate transformations, in particular affine transformations. In this appendix, we show how to efficiently construct an MPO (41) for such a transformation.
Rather than striving for full generality, we limit our discussion to the type of affine transformations needed in this paper: transformations between the native frequency representations for the ph, pp, and $\overline{\textit{ph}}$ channels. In practice, we limit the range of Matsubara frequencies to a finite box. The frequencies within that box can be enumerated by positive integers. Upon transforming to another channel, some frequencies will be mapped to lie outside the frequency box of the new channel, causing missing information in the mapping. Those frequency points then have to be dropped (open boundary conditions) or periodically continued (periodic boundary conditions). With open boundary conditions, the mapping will generically become non-invertible. For the remaining frequencies, the channel transformation maps one constrained set of positive integers to another.
We formalize the above scenario as follows. Let $\bm{x}$ and $\bm{y}$ be vectors with $N$ components. An affine transformation is a map $\bm{x}\mapsto\bm{y}$ that can be represented as
$\bm{y}=A\bm{x}+\bm{b}\,,$ (36)
where $A$ is an invertible $N\times N$ matrix. In the following, we limit our description to the case relevant for channel transforms, where all components of $\bm{x},\bm{y},\bm{b},A$ are integers. We further constrain ourselves to the case where $A^{-1}$ has integer components and the components of $\bm{b}$ are nonnegative. Given a function $g(\bm{y})$, we construct a new function $f(\bm{x})$ by a coordinate transformation
$f(\bm{x}):=g(\bm{y}(\bm{x}))\,.$ (37)
We call this type of transformation a passive affine transformation, where for a given $\bm{x}$, we define the value of the new function $f(\bm{x})$ by picking the value of the old function $g(\bm{x})$ at the transformed point $\bm{y}$. In practice, we limit $\bm{x}$, $\bm{y}$ and $\bm{b}$ to a finite box $S=\{0,\ldots,2^{R}-1\}^{N}$. Then, some $\bm{x}$ may be transformed to a $\bm{y}$ outside the box, in which case the choice of periodic or open boundary conditions becomes relevant. With periodic boundary conditions, we interpret Eq. (36) as $\bm{y}\equiv A\bm{x}+\bm{b}\pmod{2^{R}}$, where $\pmod{2^{R}}$ is to be understood component-wise. With open boundary conditions, we set $f(\bm{x})=0$ if $\bm{y}\notin S$. The transformation (36) is not necessarily invertible on $S$, even if it is invertible on $\mathbb{Z}^{N}$.
We can write Eq. (37) as a tensor product
$f(\bm{x})=\sum_{\bm{y}\in S}T(\bm{x},\bm{y})g(\bm{y})\,,$ (38)
with
$T(\bm{x},\bm{y}):=\begin{cases}1&\bm{y}=A\bm{x}+\bm{b}\,,\\ 0&\text{else.}\end{cases}$ (39)
In quantics representation, the tensor $T$ can be factorized to an MPO with small bond dimension, which allows cheap transformation of functions given in QTT format through a single MPO-MPS contraction.
To construct this MPO, it is useful to start with fused rather than interleaved indices, i.e., we only separate out the length scales of the vector, but not its components:
@@
where $\bm{x}_{r}$ is a vector of $N$ bits corresponding to the current scale, i.e., $\bm{x}_{r}\in\{0,1\}^{N}$. Thus, the legs $\bm{x}_{r}$ of the corresponding MPS are of dimension $2^{N}$ rather than $2$. We perform similar decompositions for $\bm{y}$ and $\bm{b}$. Consequentially, $T$ is decomposed as
$T(\bm{x},\bm{y})=\sum_{\alpha_{1}=1}^{D_{1}}\cdots\!\!\sum_{\alpha_{R-1}=1}^{D_{R-1}}[T_{1}]^{\bm{x}_{1}\bm{y}_{1}}_{1\alpha_{1}}[T_{2}]^{\bm{x}_{2}\bm{y}_{2}}_{\alpha_{1}\alpha_{2}}\cdots[T_{R}]^{\bm{x}_{R}\bm{y}_{R}}_{\alpha_{R-1}1},$ (41)
where $[T_{r}]^{\bm{x}_{r}\bm{y}_{r}}_{\alpha_{r-1}\alpha_{r}}$ is the $r$-th core tensor with virtual indices $\alpha_{r-1}$ and $\alpha_{r}$ as well as local indices $\bm{x}_{r}$ and $\bm{y}_{r}$. The corresponding tensor network diagram is shown in Fig. 14. The indices are bound by $\alpha_{r}\in\{1,\ldots,D_{r}\}$, $\bm{x}_{r},\bm{y}_{r}\in\{0,1\}^{N}$. Once the MPO is constructed in this way, we can transform it to the interleaved representation by splitting the core tensors using a $QR$ decomposition.
Figure 14: Affine transform $T(\bm{y},\bm{x})$ applied to function $g(\bm{x})$ in MPO form.
For the explicit construction of the MPO, we first decompose Eq. (36) for the finest scale $r=R$:
$2\bm{c}_{R-1}+\bm{y}_{R}=A\bm{x}_{R}+\bm{b}_{R}\,,$ (42)
where $\bm{c}_{R-1}$ is the carry, a vector of integers not confined to 0 and 1. Since all components of $2\bm{c}_{R-1}$ are even, we find that legal values of $\bm{y}_{R}$ must satisfy:
$\bm{y}_{R}$ $\equiv A\bm{x}_{R}+\bm{b}_{R}\pmod{2},$ (43a)
where (mod $(\mathrm{mod}\ 2)$ is to be understood component-wise. Consequently, the carry is obtained as
$\bm{c}_{R-1}$ $=\tfrac{1}{2}(A\bm{x}_{R}+\bm{b}_{R}-\bm{y}_{R}).$ (43b)
@@
The carry $\bm{c}_{R-1}$ enters the calculation for the next scale $R-1$, so we have to “communicate” it to the previous core tensor $T_{R-1}$ via the bond. To do so, we first observe that Eqs. (43) uniquely determine $\bm{y}_{R}$ and $\bm{c}_{R}$ for each $\bm{x}_{R}$. We collect all distinct values of the carry for all possible inputs $\bm{x}_{R}$ into a tuple $(\bm{c}_{R-1,1},\ldots,\bm{c}_{R-1,D_{R-1}})$. The core tensor is then given by
$[T_{R}]^{\bm{x}_{R}\bm{y}_{R}}_{\alpha 1}=\begin{cases}1&2\bm{c}_{R-1,\alpha}+\bm{y}_{R}=A\bm{x}_{R}+\bm{b}_{R},\\ 0&\mbox{else}.\end{cases}$ (44)
For all the other scales $r$, we must add the incoming carry $\bm{c}_{r}$ and must thus amend Eq. (42) to
$2\bm{c}_{r-1}+\bm{y}_{r}=A\bm{x}_{r}+\bm{b}_{r}+\bm{c}_{r}\,,$ (45)
and Eqs. (43) to:
$\bm{y}_{r}$ $\equiv A\bm{x}_{r}+\bm{b}_{r}+\bm{c}_{r}\pmod{2}\,,$ (46a)
$\bm{c}_{r-1}$ $=\tfrac{1}{2}(A\bm{x}_{r}+\bm{b}_{r}+\bm{c}_{r}-\bm{y}_{r}).$ (46b)
@@
We again collect all distinct outgoing carry values for all possible $\bm{x}_{r}$ and $\bm{c}_{r}$ into $(\bm{c}_{r-1,\alpha})_{\alpha=1,\ldots,D_{r-1}}$, and obtain the core tensor
$[T_{r}]^{\bm{x}_{r}\bm{y}_{r}}_{\alpha\alpha^{\prime}}=\begin{cases}1&2\bm{c}_{r-1,\alpha}+\bm{y}_{r}=A\bm{x}_{r}+\bm{b}_{r}+\bm{c}_{r,\alpha^{\prime}},\\ 0&\mbox{else}.\end{cases}$ (47)
We iterate this procedure from $r=R$ to $1$, constructing all MPO core tensors in a single backward sweep. Having reached the first tensor, $r=1$, we implement open boundary conditions by demanding that $\bm{c}_{0}=\bm{0}$ in Eq. (47), such that
$[T_{1}]^{\bm{x}_{1}\bm{y}_{1}}_{1\alpha^{\prime}}=\begin{cases}1&\bm{y}_{1}=A\bm{x}_{1}+\bm{b}_{1}+\bm{c}_{1,\alpha^{\prime}},\\ 0&\mbox{else}.\end{cases}$ (48)
Periodic boundary conditions are implemented by modifying Eq. (47) such that the leftmost carry $\bm{c}_{0}$ is discarded, such that
@@
$[T_{1}]^{\bm{x}_{1}\bm{y}_{1}}_{1\alpha^{\prime}}=\begin{cases}1&\bm{y}_{1}\equiv A\bm{x}_{1}+\bm{b}_{1}+\bm{c}_{1,\alpha^{\prime}}\pmod{2},\\ 0&\mbox{else}.\end{cases}$ (49)
The bond dimension $D_{\mathrm{max}}=\max_{r}D_{r}$ of the tensors constructed in this way is likely optimal. This algorithm has $\mathcal{O}(RD_{\mathrm{max}}^{2}2^{2N})$ runtime, which is optimal in the sense that at least this amount of runtime and memory is necessary to construct the tensors $T_{r}$. The algorithm can be generalized to cases where $A$ and $\bm{b}$ have entries in $\mathbb{Q}$.
$r$ $\alpha_{r-1}$ $\bm{c}_{r-1}$ $\bm{y}_{r}$ $\bm{x}_{r}$ $\alpha_{r}$ $\bm{c}_{r}$
$R$ $1$ $(0,\phantom{-}0)$ $(0,0)$ $(0,0)$ $1$ $(0,\phantom{-}0)$
$1$ $(0,\phantom{-}0)$ $(1,1)$ $(1,0)$ $1$ $(0,\phantom{-}0)$
$2$ $(0,-1)$ $(0,1)$ $(0,1)$ $1$ $(0,\phantom{-}0)$
$1$ $(0,\phantom{-}0)$ $(1,0)$ $(1,1)$ $1$ $(0,\phantom{-}0)$
$R-1$ $1$ $(0,\phantom{-}0)$ $(0,0)$ $(0,0)$ $1$ $(0,\phantom{-}0)$
$1$ $(0,\phantom{-}0)$ $(1,1)$ $(1,0)$ $1$ $(0,\phantom{-}0)$
$2$ $(0,-1)$ $(0,1)$ $(0,1)$ $1$ $(0,\phantom{-}0)$
$1$ $(0,\phantom{-}0)$ $(1,0)$ $(1,1)$ $1$ $(0,\phantom{-}0)$
$2$ $(0,-1)$ $(0,1)$ $(0,0)$ $2$ $(0,-1)$
$1$ $(0,\phantom{-}0)$ $(1,0)$ $(1,0)$ $2$ $(0,-1)$
$2$ $(0,-1)$ $(0,0)$ $(0,1)$ $2$ $(0,-1)$
$2$ $(0,-1)$ $(1,1)$ $(1,1)$ $2$ $(0,-1)$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
$1$ $1$ $(0,\phantom{-}0)$ $(0,0)$ $(0,0)$ $1$ $(0,\phantom{-}0)$
$1$ $(0,\phantom{-}0)$ $(1,1)$ $(1,0)$ $1$ $(0,\phantom{-}0)$
$1$ $(0,\phantom{-}0)$ $(1,0)$ $(1,1)$ $1$ $(0,\phantom{-}0)$
$1$ $(0,\phantom{-}0)$ $(1,0)$ $(1,0)$ $2$ $(0,-1)$
Table 1: Non-zero elements of $[T_{r}]^{\bm{x}_{r}\bm{y}_{r}}_{\alpha_{r-1}\alpha_{r}}$ (41) constructed from Eqs. (42)–(48) for the affine transform (50) with open boundary conditions.
Let us walk through the algorithm for the example transformation with $N=2$ and open boundary conditions:
$\bm{y}=\begin{pmatrix}1&\phantom{-}0\\ 1&-1\end{pmatrix}\bm{x}+\begin{pmatrix}0\\ 0\end{pmatrix}.$ (50)
The non-zero elements of the corresponding core tensors (41) are listed in Table 1. For the case $r=R$, we simply apply the transformation (43a) to each bit combination $\bm{x}_{R}\in\{(0,0),(1,0),(0,1),(1,1)\}$. In the case $\bm{x}_{R}=(0,1)$, Eq. (43b) yields a carry of $\bm{c}_{R-1}=(0,-1)$, to which we assign the bond index $\alpha_{R-1}=2$, otherwise it is $(0,0)$, to which we assign the index $\alpha_{R-1}=1$. The dimension of the corresponding bond is thus $D_{R-1}=2$ and the core tensor has the four nonzero elemnts listed in rows 1–4 of Table 1.
For $r=R-1$, Eq. (46) directs us to add the incoming carry $\bm{c}_{r}$. Hence, we double the number of non-zero entries, as we have to repeat the calculation for each of the two outgoing carries of $T_{R}$. We observe that the set of incoming and outgoing carries is identical and assign the same bond indices to them. The corresponding nonzero elements of $T_{R-1}$ are then listed in rows 5–12 of Table 1. Since the values of the outgoing carries form the same set as those of the incoming carries, all other core tensors $T_{r^{\prime}}$ with $1<r^{\prime}<R$ are equal to $T_{R-1}$. For $r=1$, we impose open boundary conditions, thereby restricting the outgoing carry of $T_{1}$ to zero, as shown in Eq. (48). This cuts half of the elements and yields the entries listed in rows 13–16 of Table 1.